### v1

# Source-Free Domain Adaptation with Diffusion-Guided Source Data Generation

Abstract
This paper introduces a novel approach to leverage the generalizability capability of Diffusion Models for Source-Free Domain Adaptation (DM-SFDA). Our proposed DM-SFDA method involves fine-tuning a pre-trained text-to-image diffusion model to generate source domain images using features from the target images to guide the diffusion process. Specifically, the pre-trained diffusion model is fine-tuned to generate source samples that minimize entropy and maximize confidence for the pre-trained source model. We then apply established unsupervised domain adaptation techniques to align the generated source images with target domain data. We validate our approach through comprehensive experiments across a range of datasets, including Office-31 (Saenko et al., 2010), Office-Home (Venkateswara et al., 2017), and VisDA (Peng et al., 2017). The results highlight significant improvements in SFDA performance, showcasing the potential of diffusion models in generating contextually relevant, domain-specific images.
1 Introduction
Deep Convolutional Neural Networks (CNNs) have demonstrated impressive performance on several visual tasks in recent years. However, the assumption that the distributions of the training and test sets are the same is crucial to the effectiveness of CNNs (Ding et al., 2022). Subsequently, a big drop in performance is generally observed when CNN-based models are deployed in real-world settings with a discrepancy in data distribution (Kundu et al., 2020). Domain Adaptation (DA) attempts to reduce this disparity to make these models perform well across multiple domains. Traditional DA approaches that rely on fixed source data might struggle to keep up with the pace of domain changes. Moreover, the rising prominence of data privacy regulations has led to a demand for DA techniques that can function effectively without relying on access to the source training data, a setting generally known as Source Free Domain Adaptation (SFDA).
Most of the current state-of-the-art DA methods attain model adaptability by converging two disparate data distributions within a shared feature space, spanning both domains simultaneously (Ding et al., 2022). One way of achieving this in a source-free manner is to use synthetically generated source data. However, generating synthetic source data that accurately represents the diversity and complexity of the source domain can be difficult. Furthermore, if the synthetic data is not of high quality, it might introduce noise and inconsistencies, negatively impacting the model’s performance on the target domain. Notably, recent advancements in Diffusion Generative Models (DGMs) (Ho et al., 2020; Sohl-Dickstein et al., 2015) have demonstrated exceptional capabilities in producing diverse and high-quality images. Consequently, this paper aims to harness the generalizability of the state-of-the-art text-to-image diffusion models to the challenging task of SFDA.
To address the challenges of data privacy and diversity in the reconstruction process, we present an innovative framework named Diffusion Models for Source-Free Domain Adaptation (DM-SFDA). An overview of DM-SFDA is illustrated in Figure 1. The core idea of this approach is to use text-to-image diffusion models to generate images representative of the source domain based on the target domain and a pre-trained source network. Essentially, this involves fine-tuning a pre-trained text-to-image diffusion model to produce source samples that minimize the entropy for the pre-trained source model. The key contributions of our framework can be summarized as follows.
The remainder of the paper is organized as follows. In Section 2, we discuss related work in the areas of DA, SFDA, and DGMs. In Section 3, we revisit the preliminary concepts that form the basis of our proposed approach. In Section 4, we discuss and formalize the problem definition and present our methodology. We present experimental results and analysis in Sections 5, 6, and 7. Finally, we conclude with a brief summary in Section 8.
2 Related Work
2.1 Domain Adaptation
DA has its roots in (Ben-David et al., 2006), which focused on the role of good feature representations in their successful application for the task. The initial few works in DA adapted moment matching to align feature distributions between source and target domains (Long et al., 2015b; Huang et al., 2006; Long et al., 2016a; Sun et al., 2015; Tzeng et al., 2014). Subsequent works used adversarial learning-based approaches to tackle the problem of DA (Ganin & Lempitsky, 2014; Saito et al., 2017; Long et al., 2017; Ganin et al., 2016; Shu et al., 2018). Apart from this, many other techniques like (Xu et al., 2018; Tang et al., 2020a; Saito et al., 2020; Bermúdez-Chacón et al., 2020) have been proposed to tackle the task of DA.
2.2 Unsupervised Domain Adaptation
Unsupervised Domain Adaptation (UDA) is a subtype of DA that aims to transfer knowledge from a labeled source domain to a different unlabeled target domain (Long et al., 2016b). Existing mainstream UDA methods can be categorized into two main types of methods: those that align the source and target domain distributions by designing specific metrics (Na et al., 2022, 2021; Li et al., 2021; Long et al., 2015a), and those that learn domain-invariant feature representations through adversarial learning (Westfechtel et al., 2024; Ganin & Lempitsky, 2015b; Long et al., 2018). However, the success of most of these methods depends on a huge amount of source data which might not be available in most practical scenarios.
2.3 Source-Free Domain Adaptation
SFDA has been considered in the literature as a means of reducing reliance on source data. As described in (Yu et al., 2023), the existing SFDA research can generally be categorized into two approaches: data-centric and model-centric. Model-centric methods employ techniques such as self-training and self-attention, while data-centric methods include domain-based reconstruction and image-based information extraction. Our proposed method follows the data-centric perspective to solve the SFDA task using source domain generation. 3C-GAN (Li et al., 2020) is a pioneering work in this area which uses a Generative Adversarial Network (GAN) to generate source-like images. Some other works like SDDA (Kurmi et al., 2021) and CPGA (Qiu et al., 2021) also solve the SFDA task using a similar approach.
2.4 DGMs for Domain Adaptation
Recently, there has been a significant shift in the landscape of generative modeling due to DGMs (Ho et al., 2020; Sohl-Dickstein et al., 2015), demonstrating impressive capabilities in generating highly realistic text-conditioned images. DGMs have also seen a growing interest in the DA community with many recent works using DGMs as input augmentation techniques. A recent example is a text-to-image diffusion model, employed by (Benigmim et al., 2023) to generate target domain images using source domain labels, thereby demonstrating the efficacy of diffusion models in One-Shot Unsupervised Domain Adaptation (OSUDA). DGMs, when trained on multiple source domains, have also been instrumental in guiding approximate inference in target domains, as reported by (Graikos et al., 2022). In our work, we leverage a recently introduced fine-tuning strategy for diffusion models called Denoising Diffusion Policy Optimization (DDPO) (Black et al., 2023), to fine-tune the diffusion models using the output probability of the source model as an objective function.
3 Preliminaries
3.1 Conditional Diffusion Probabilistic Models
Conditional Denoising Diffusion Probabilistic Model (DDPM) forms the backbone of our source data reconstruction pipeline. This model represent a distribution over data , conditioned on a contextual input labeled . This distribution arises from a sequential denoising process, which aims to reverse a Markovian forward process denoted as . This forward process progressively introduces a Gaussian noise to the data. Subsequently, a forward process posterior mean predictor is then trained to reverse the forward process for all . The training process entails maximizing a variational lower bound on the model log-likelihood with an objective function defined as follows:
where is a weighted average of and .
The sampling process of a diffusion model starts with a sample from the Gaussian distribution which is subsequently denoised using the reverse process to product a trajectory ending with a sample . Here, the sampler use an isotropic Gaussian reverse process with a fixed time-dependent variance:
3.2 Markov Decision Process and Reinforcement Learning
A Markov decision process (MDP) serves as a structured representation of problems involving a series of interconnected decisions. It is characterized by a set of components denoted as , where signifies the collection of possible states, stands for the available actions, signifies the initial state distribution, represents the transition pattern, and defines the reward mechanism. At each discrete time step denoted as , the agent observes a specific state represented as from the state space takes an action labeled as from the action space and in response receives a reward labeled as , subsequently transitioning to a novel state drawn from the distribution . The agent’s decision-making is guided by a policy that dictates actions based on states. During the agent’s engagement with the MDP, it generates trajectories, which are sequences comprising both states and actions, conventionally presented as . Within the context of reinforcement learning (RL), the primary aim of the agent is to maximize a metric termed , representing the expected cumulative reward garnered over trajectories sampled in accordance with its policy:
3.3 Denoising as a multi-step MDP
When the sampler is fixed, the denoising process can be reframed as a multi-step MDP (Black et al., 2023). The iterative denoising process can be mapped to the following MDP:
Here, is the Dirac delta distribution with a non-zero density only at .
4 Proposed Method
4.1 Notations
Following the notations used in (Yu et al., 2023), in this paper, we represent a domain as . Each domain consists of a dataset and an associated label set . A dataset comprises of an instance set , derived from a -dimensional marginal distribution , and a label set where represents the total number of samples and represents the total number of classes.
4.2 Problem Definition
The SFDA scenario involves two stages: pre-training and adaptation. During pre-training, a model is trained on labeled data from the source domain . Subsequently, the goal of the adaptation phase is to adapt the pre-trained source model to the unlabeled target data . The proposed approach assumes a closed form, implying that the label spaces of the source and target domains are identical.
4.3 Method Overview
An overview of the proposed method DM-SFDA for source free domain adaptation using text-to-image diffusion models can be seen in Figure 1. The key idea behind the approach is to leverage the generalizability of the state-of-the-art text-to-image diffusion models to tackle the task of SFDA. As shown in Figure 2, the training pipeline contains the following four phases: I) Selective Pseudo Labeling Target Data II) Fine-tuning Diffusion Model on Target Data, III) Source Data Generation using DDPO, IV) Unsupervised Domain Adaptation. The first phase involves selective pseudo labeling target data using the pre-trained source model. The second phase involves jointly fine-tuning an implicit image captioner along with a pre-trained text-to-image diffusion model on the target images. Subsequently, the pre-trained source model is used to fine-tune this diffusion model using Denoising Diffusion Policy Optimization (DDPO) (Black et al., 2023) to generate Source Images. Finally, the generated source images and the target images are used to perform unsupervised domain adaptation. Each of these phases are described in detail in the following subsections.
4.4 Phase I: Selective Pseudo Labeling Target Data
The proposed DM-SFDA pipeline requires labeled target data to generate data from the source domain. Therefore, the initial phase of our proposed pipeline addresses the challenge of unlabeled target data by selecting reliable labels for the target samples using a selective pseudo-labeling strategy. This approach is akin to the one proposed in (Karim et al., 2023). As shown in (Guillory et al., 2021), prediction confidence and difference in entropy can be reliable measures of pseudo-label accuracy and estimate different types of domain shifts. Therefore, prediction confidence and average entropy are used as metrics to assess label reliability. A binary reliability score () is assigned to each sample in the target data, determined by their prediction confidence and prediction uncertainty (), as illustrated below
Here, and represent the selection thresholds for confidence and uncertainity. Taking the average as a threshold eliminates the requirement of per-dataset hyper-parameter tuning and makes the selection process highly adaptive. Furthermore, aleatoric uncertainty (Hüllermeier & Waegeman, 2021) is used since it better addresses the concern of domain shift.
4.5 Phase II: Diffusion Model Finetuning on Target Data
The second phase of the proposed DM-SFDA pipeline involves fine-tuning a text-to-image diffusion model on the target data. However, the unavailability of labels classes poses a challenge as we don’t have any textual cues to guide the diffusion process. Furthermore, passing an empty caption would also be sub-optimal (Xu et al., 2023). This issue of unavailability of target labels is tackled by using an implicit image captioner, as used in (Xu et al., 2023). Instead of using an off-the-shelf captioning network to generate captions, a network is trained to generate an implicit text embedding from the input image itself. The orange block in Figure 2 shows the architecture of the implicit image captioner. Specifically, to derive the implicit text embedding of the image, a pre-trained frozen image encoder, (e.g. from CLIP (Radford et al., 2021) is used to encode the input image into its embedding space. A learned MLP is added to project the image embedding into an implicit text embedding. The image captioner is jointly finetuned along with the diffusion model to generate implicit text embedding for the input target image. During the training phase, the parameters of the image encoder is unchanged and we only finetune the parameters of the MLP and UNet. Formally, the textual features of a target image are extracted as follows:
4.6 Phase III: Source Data Generation using DDPO
In the third phase, we use the method proposed in (Black et al., 2023) to further fine-tune our diffusion model to generate source-like images. As explained in (Black et al., 2023), when we formulate the denoising process of a diffusion model as a multi-step MDP, and we use a standard sampler parameterized as in Equation 2, the policy becomes an isotropic Gaussian. This simplification allows for the evaluation of exact action likelihoods and gradients of these likelihoods with respect to the diffusion model parameters. Subsequently, with access to likelihood and likelihood gradients, the fine-tuning alternates collecting trajectories via sampling and updating parameters via gradient ascent on the diffusion RL objective :
The pre-trained source model serves as the reward function for the fine-tuning procedure. The ultimate objective is to maximize the average classification confidence and minimize the average entropy on the generated source data. The reward function used for fine-tuning the model is shown below:
After the completion of the DDPO fine-tuning, the diffusion model produces source images by utilizing the implicit target image captions as guiding prompts.
4.7 Phase IV: Unsupervised Domain Adaptation
The fourth phase starts by labeling the generated source domain images using the pre-trained source model. Once we have the reconstructed and labeled source domain data, we have effectively converted the initial problem of SFDA to a standard Unsupervised Domain Adaptation (UDA) problem which can be solved using any of the pre-existing approaches. Since we chose a ResNet backbone as the pre-trained source model, we experiment with the current state-of-the-art ResNet-based UDA approaches. For the purpose of this paper, we experiment with the following approaches:
4.7.1 Bridging Domain Spaces (FixBi)
FixBi (Na et al., 2021) tackles the problem of domain adaptation by compensating for the large discrepancy between the source and target domains. This is done by generating multiple intermediate augmented domains by using a fix ration-based mixup data augmentation approach. Complementary models are trained on each of these intermediate domains that teach each other to bridge the domain gap by using a confidence-based learning where one model teaches the other model using the positive pseudo-labels or teach itself using the negative pseudo-labels. Through the confidence-based learning approach, the two models with different characteristics gradually get closer to the target domain. Furthermore, a consistency regularization is used to ensure a stable convergence of training both models.
| Method |
Avg. | ||||||||
| ResNet-50 (He et al., 2015) | ✗ | 68.9 | 68.4 | 62.5 | 96.7 | 60.7 | 99.3 | 76.1 | ||
| MCC (Jin et al., 2019) | ✗ | 95.6 | 95.4 | 72.6 | 98.6 | 73.9 | 100.0 | 89.4 | ||
| GSDA (Hu et al., 2020) | ✗ | 94.8 | 95.7 | 73.5 | 99.1 | 74.9 | 100.0 | 89.7 | ||
| SRDC (Tang et al., 2020b) | ✗ | 95.8 | 95.7 | 76.7 | 99.2 | 77.1 | 100.0 | 90.8 | ||
| FixBi (Na et al., 2021) | ✗ | 95.0 | 96.1 | 78.7 | 99.3 | 79.4 | 100.0 | 91.4 | ||
| CoVi (Na et al., 2022) | ✗ | 98.0 | 97.6 | 77.5 | 99.3 | 78.4 | 100.0 | 91.8 | ||
| SHOT (Liang et al., 2020) | ✓ | 94.0 | 90.1 | 74.7 | 98.4 | 74.3 | 99.9 | 88.6 | ||
| 3C-GAN (Li et al., 2020) | ✓ | 92.7 | 93.7 | 75.3 | 98.5 | 77.8 | 99.8 | 89.6 | ||
| NRC (Yang et al., 2021) | ✓ | 96.0 | 90.8 | 75.3 | 99.0 | 75.0 | 100.0 | 89.4 | ||
| NRC++ (Yang et al., 2023) | ✓ | 95.9 | 91.2 | 75.5 | 99.1 | 75.0 | 100.0 | 89.5 | ||
| DM-SFDA (FixBi) | ✓ | 93.0 | 93.5 | 77.4 | 98.7 | 78.3 | 99.7 | 90.1 | ||
| DM-SFDA (CoVi) | ✓ | 93.6 | 94.0 | 77.0 | 99.0 | 78.0 | 99.9 | 90.2 |
| Method | Source | Ar | Cl | Pr | Rw | Avg. | ||||||||
| Free | Cl | Pr | Rw | Ar | Pr | Rw | Ar | Cl | Rw | Ar | Cl | Pr | ||
| ResNet-50 (He et al., 2015) | ✗ | 34.9 | 50.0 | 58.0 | 37.4 | 41.9 | 46.2 | 38.5 | 31.2 | 60.4 | 53.9 | 41.2 | 59.9 | 46.1 |
| MCC (Jin et al., 2019) | ✗ | 88.1 | 80.3 | 80.5 | 71.5 | 90.1 | 93.2 | 85.0 | 71.6 | 89.4 | 73.8 | 85.0 | 36.9 | 78.8 |
| GSDA (Hu et al., 2020) | ✗ | 61.3 | 76.1 | 79.4 | 65.4 | 73.3 | 74.3 | 65.0 | 53.2 | 80.0 | 72.2 | 60.6 | 83.1 | 70.3 |
| SRDC (Tang et al., 2020b) | ✗ | 52.3 | 76.3 | 81.0 | 69.5 | 76.2 | 78.0 | 68.7 | 53.8 | 81.7 | 76.3 | 57.1 | 85.0 | 71.3 |
| FixBi (Na et al., 2021) | ✗ | 58.1 | 77.3 | 80.4 | 67.7 | 79.5 | 78.1 | 65.8 | 57.9 | 81.7 | 76.4 | 62.9 | 86.7 | 72.7 |
| CoVi (Na et al., 2022) | ✗ | 58.5 | 78.1 | 80.0 | 68.1 | 80.0 | 77.0 | 66.4 | 60.2 | 82.1 | 76.6 | 63.6 | 86.5 | 73.1 |
| SHOT (Liang et al., 2020) | ✓ | 57.1 | 78.1 | 81.5 | 68.0 | 78.2 | 78.1 | 67.4 | 54.9 | 82.2 | 73.3 | 58.8 | 84.3 | 71.8 |
| NRC (Yang et al., 2021) | ✓ | 57.7 | 80.3 | 82.0 | 68.1 | 79.8 | 78.6 | 65.3 | 56.4 | 83.0 | 71.0 | 58.6 | 85.6 | 72.2 |
| NRC++ (Yang et al., 2023) | ✓ | 57.8 | 80.4 | 81.6 | 69.0 | 80.3 | 79.5 | 65.6 | 57.0 | 83.2 | 72.3 | 59.6 | 85.7 | 72.5 |
| DM-SFDA (FixBi) | ✓ | 57.9 | 77.1 | 80.1 | 67.3 | 79.4 | 78.0 | 65.7 | 57.5 | 81.6 | 75.3 | 62.7 | 86.5 | 72.4 |
| DM-SFDA (CoVi) | ✓ | 58.0 | 77.8 | 79.6 | 67.9 | 80.0 | 76.8 | 66.0 | 58.8 | 82.0 | 75.8 | 63.0 | 86.3 | 72.6 |
| Method | Source-Free | plane | bike | bus | car | horse | knife | mcycle | person | plant | sktbrd | train | truck | Avg. |
| ResNet-101 (He et al., 2015) | ✗ | 55.1 | 53.3 | 61.9 | 59.1 | 80.6 | 17.9 | 79.7 | 31.2 | 81.0 | 26.5 | 73.5 | 8.5 | 52.4 |
| MCC (Jin et al., 2019) | ✗ | 88.7 | 80.3 | 80.5 | 71.5 | 90.1 | 93.2 | 85.0 | 71.6 | 89.4 | 73.8 | 85.0 | 36.9 | 78.8 |
| GSDA (Hu et al., 2020) | ✗ | 93.1 | 67.8 | 83.1 | 83.4 | 94.7 | 93.4 | 93.4 | 79.5 | 93.0 | 88.8 | 83.4 | 36.7 | 81.5 |
| FixBi (Na et al., 2021) | ✗ | 96.1 | 87.8 | 90.5 | 90.3 | 96.8 | 95.3 | 92.8 | 88.7 | 97.2 | 94.2 | 90.9 | 25.7 | 87.2 |
| CoVi (Na et al., 2022) | ✗ | 96.8 | 85.6 | 88.9 | 88.6 | 97.8 | 93.4 | 91.9 | 87.6 | 96.0 | 93.8 | 93.6 | 48.1 | 88.5 |
| SHOT (Liang et al., 2020) | ✓ | 94.3 | 88.5 | 80.1 | 57.3 | 93.1 | 94.9 | 80.7 | 80.3 | 91.5 | 89.1 | 86.3 | 58.2 | 82.9 |
| 3C-GAN (Li et al., 2020) | ✓ | 94.8 | 73.4 | 68.8 | 74.8 | 93.1 | 95.4 | 88.6 | 84.7 | 89.1 | 84.7 | 83.5 | 48.1 | 81.6 |
| NRC (Yang et al., 2021) | ✓ | 96.8 | 91.3 | 82.4 | 62.4 | 96.2 | 95.9 | 86.1 | 80.6 | 94.8 | 94.1 | 90.4 | 59.7 | 85.9 |
| NRC++ (Yang et al., 2023) | ✓ | 96.8 | 91.9 | 88.2 | 82.8 | 97.1 | 96.2 | 90.0 | 81.1 | 95.2 | 93.8 | 91.1 | 49.6 | 87.8 |
| DM-SFDA (FixBi) | ✓ | 95.9 | 87.6 | 89.0 | 86.2 | 96.6 | 95.0 | 91.3 | 86.4 | 96.5 | 94.0 | 90.4 | 30.2 | 86.6 |
| DM-SFDA (CoVi) | ✓ | 96.3 | 86.0 | 88.7 | 85.4 | 97.1 | 93.4 | 91.0 | 86.0 | 95.8 | 93.7 | 91.3 | 46.6 | 87.5 |
4.7.2 Contrastive Viscinal Space (CoVi)
CoVi (Na et al., 2022) utilises the viscinal space between the source and target domain for unsupervised domain adaptation. A Contrastive Viscinal Space-based algorithm is introduced, where the viscinal space is divided into contrastive space (source dominant) and consensus space (target dominant) by using the Entropy Maximization Point (EMP) as a boundary to resolve the problem of equilibrium collapse of labels. In the contrastive space, the inter-domain discrepancy is mitigated by constraining instances to have contrastive views and labels. The consensus space reduces the confusion between intra-domain categories by enforcing label consensus among the multiple target-dominant vicinal instances to a single target label.
5 Experiments and Results
5.1 Datasets
Office-31: Office-31 (Saenko et al., 2010) is a benchmark image classification dataset that consists of a limited set of images distributed across 31 categories spanning three domains: Amazon (2,817 images), DSLR (498 images), and Webcam (795 images).
Office-Home: Office-Home (Venkateswara et al., 2017), on the other hand, comprises a more extensive dataset with a total of 15.5K images from 65 classes, gathered from 4 distinct image domains: Artistic, Clipart, Product, and Real-world. Our analysis includes 12 transfer tasks for this dataset.
VisDA: VisDA (Peng et al., 2017) encompasses two distinct domains: synthetic and real, each comprising 12 classes. The synthetic domain holds around 150K computer-generated 3D images with various poses, while the corresponding real domain includes approximately 55K images captured from the real world.
5.2 Evaluation Metric
Our evaluation metrics encompass per-class accuracy and their average (Avg.) performance. Our evaluation involves transfer tasks spanning all the domains of each dataset, and we report the top-1 accuracy achieved by various methods under different settings.
5.3 Experimental Setup
We implement our approach using PyTorch and use ResNet-50 (He et al., 2015) as the backbone network for the Office-31 (Saenko et al., 2010) and Office-Home (Venkateswara et al., 2017) datasets and Resnet-101 for the VisDA (Peng et al., 2017) dataset. All the experiments for our proposed approach were conducted on a Nvidia A100 GPU. The other details for specific parts of our pipeline are specified in the subsequent sub-sections.
5.3.1 Diffusion Model Fine-tuning
We use Stable Diffusion v1.4 (Rombach et al., 2022) as the base model for all experiments. The finetuning of the diffusion models was done in the Accelerate (Gugger et al., 2022) environment. Memory-efficient attention was enabled using xFormers (Lefaudeux et al., 2022) for all the experiments performed. Furthermore, we used the Low-Rank Adaptation (LORA) of the Stable Diffusion Pipeline.
5.3.2 Denoising Diffusion Policy Optimization
In the context of the DDPO experiments, we employed the Low-Rank Adaptation (LORA) technique within the framework of the Stable Diffusion Pipeline. The training was done for a batch size of 4, and 100 batches were sampled per step. The training procedure encompassed 100 steps, with each step incorporating two distinct phases: a sampling phase and 10 consecutive inner training epochs dedicated to training the model on the sampled data from the previous phase.
5.3.3 Unupervised Domain Adaptation
In the unsupervised domain adaptation phase of our pipeline, we use the experimental setup proposed in each individual approaches (Na et al., 2021, 2022). For the UDA approaches, we use the generated source data labeled using the pre-trained model and all unlabeled target data. We use minibatch stochastic gradient descent (SGD) with a momentum of 0.9, an initial learning rate of 0.001, and a weight decay of 0.005. We follow the same learning rate schedule as in (Ganin & Lempitsky, 2015a).
5.4 Results
5.4.1 Results on Office-31
We present the outcomes for the Office-31 dataset in Table 1. By employing the source data produced through our proposed pipeline as input, the UDA methods managed to surpass the existing state-of-the-art SFDA approaches by margins of 0.6% and 0.7%, achieving average accuracies of 90.1% and 90.2% with the FixBi (Na et al., 2021) and CoVi (Na et al., 2022) approaches respectively.
5.4.2 Results on Office-Home
The results for the Office-Home dataset (Venkateswara et al., 2017) are presented in Table 2. By integrating the source data generated through our proposed pipeline into the UDA methodologies, we have successfully surpassed existing methods, achieving an average accuracy of 72.4% and 72.6% with FixBi (Na et al., 2021) and CoVi (Na et al., 2022) respectively.
5.4.3 Results on VisDA17
We summarize the results for the VisDA dataset (Peng et al., 2017) in Table 3. Our proposed framework is able to outperform the source-free baselines across many classes like bus and car by 0.8% and 3.4% respectively. The accuracies of our proposed method are very close to the approaches where source data is available, thereby showing the effectiveness of our data generation pipeline.
6 Ablation Study
In this section, we perform an ablation study to understand the contribution of each component in the proposed pipeline to the overall performance.
6.1 Selective Pseudo Labeling Target Data
Correct pseudo labeling of target data samples is essential for the success of the proposed approach, as it significantly influences the fine-tuning process of the diffusion model. To gauge its impact, we assess the model’s performance without employing selective pseudo labeling. This will help assess the importance of initial pseudo labels in guiding the subsequent fine-tuning and adaptation phases. As shown in Table 4, a significant drop in performance is observed for the proposed pipeline in the absence of the selective pseudo labeling. The primary reason for this is the inaccuracies in pseudo label assignment during the initial phase that adversely affect all subsequent phases of the pipeline, including data generation and UDA.
Avg. | ||||||||
| ✗ | 67.8 | 68.3 | 60.1 | 95.4 | 60.5 | 98.7 | 75.13 | ||
| ✓ | 93.6 | 94.0 | 77.0 | 99.0 | 78.0 | 99.9 | 90.2 |
6.2 Image Caption Generators
The output of a text-to-image diffusion model is conditioned on the input prompt. To derive the optimal output from the diffusion model, we used an implicit captioner as described in Section 4.4. In this context, we explored the use of an off-the-shelf Vision Language Model, LlaVa (Liu et al., 2023), to produce captions from images, which then served as inputs for the diffusion models. The outcomes from the caption generator ablation study are consolidated in Table 5. Our findings indicate that fine-tuning an implicit captioner yields superior results compared to utilizing an off-the-shelf model.
Avg. | |||||||
| LLaVa | 90.5 | 92.0 | 75.2 | 95.4 | 76.1 | 99.5 | 88.1 | |
| Implicit Captioner | 93.6 | 94.0 | 77.0 | 99.0 | 78.0 | 99.9 | 90.2 |
7 Discussion
7.1 Visualization Analysis
Figure LABEL:fig:tsne illustrates the t-SNE visualizations (van der Maaten & Hinton, 2008) for the latent feature space of the target domain, post-adaptation, using the Office-31 dataset (Saenko et al., 2010). It is evident that the target instances are significantly clustered alongside the source instances within each class following the adaptation process. This clustering underscores the model’s capability to perform accurate class conditional generation within the target domain, thereby showcasing its superior performance in data generation. To further demonstrate the effectiveness of our proposed approach, we visually inspect the generated source domain images. Figure 3 shows the data generated for the Office-31 dataset for the DSLR Amazon task. As shown, the proposed approach was able to successfully generate images similar to the source domain.
7.2 Challenges and Limitations
Computational Resources: Training and running the proposed pipeline to generate source-like images is computationally intensive and time-consuming, requiring significant computational resources. This may limit the practicality of the proposed approach for researchers or practitioners with limited access to such resources.
Scalability to Different Domains: The effectiveness of the proposed method in different application domains has yet to be fully explored. Some domains might present unique challenges that are not adequately addressed by the current model, such as highly structured where slight inaccuracies in generated data could lead to significant performance drops.
8 Conclusion
In this work, we presented a novel approach to tackle the problem of SFDA by using a text-to-image diffusion model to generate source images with high confidence predictions by the pre-trained models. We conduct extensive experiments on multiple domain adaptation benchmarks. Compared with recent data-based domain adaptation methods, our model achieves the best or comparable performance in the absence of source data, thereby proving the efficacy of our proposed approach.
9 Impact Statement
The proposed work introduces a novel approach for SFDA through the regeneration of source data using a text-to-image diffusion model. The implications of this work span several dimensions of societal, ethical, and technological realms, promising to influence various sectors positively while also presenting areas requiring careful consideration.
9.1 Societal Impact
Our methodology has the potential to democratize access to advanced machine learning capabilities by reducing the dependency on large, annotated datasets. By enabling effective domain adaptation without direct access to source data, we can foster innovation in areas where data privacy concerns or proprietary restrictions limit data availability. Applications in healthcare, environmental monitoring, and autonomous systems could see accelerated advancements, improving services and outcomes in these critical sectors.
9.2 Ethical Considerations
The use of pre-trained models and generative techniques raises important ethical questions. While our approach aims to mitigate data availability challenges, it also emphasizes the need for vigilance against perpetuating biases present in pre-trained models. Generated images must be scrutinized for biases that could lead to unfair or prejudiced outcomes, especially in sensitive applications. Moreover, the ethical use of generative models must be guided by principles that prevent misuse, such as generating misleading or harmful content.
9.3 Technological Advancements
Beyond immediate applications, this work contributes to the broader field of machine learning by exploring novel uses of diffusion models in domain adaptation. It challenges current paradigms and opens new avenues for research in unsupervised learning, model generalization, and the efficient utilization of pre-trained models. As the technology matures, it could lead to more robust, adaptable systems capable of learning from diverse and limited data sources.